OpenStudy (anonymous):

Describe the end behavior of the graph for each function. 1) f(x)=x^3+2x^2-5x+1 2) g(x)=-2x^3-8x^2+18x+72

OpenStudy (dessyj1):

Hi.

OpenStudy (dessyj1):

Do you remember the correlation between the highest degree of a polynomial and its end behaviour?

OpenStudy (anonymous):

Yep! i'm just not sure how to properly describe it

OpenStudy (anonymous):

would i just say down on the left and up on the right? is that correct for 1?

OpenStudy (dessyj1):

Yes

OpenStudy (anonymous):

could i use the infinity signs or would that not work for this??

OpenStudy (dessyj1):

Or the graph is going negative when going to the left, and positive when going to the right.

OpenStudy (dessyj1):

You can use infinity signs. Infinity signs would actually be more accurate.

OpenStudy (anonymous):

how would i write that?

OpenStudy (dessyj1):

That really depends.

OpenStudy (dessyj1):

@ganeshie8

OpenStudy (dessyj1):

I would use limits

OpenStudy (anonymous):

I don't get it :/

OpenStudy (dessyj1):

I will show you what i mean.

OpenStudy (dessyj1):

\[\lim_{x \rightarrow -\infty} = -\infty\]

OpenStudy (dessyj1):

\[\lim_{x \rightarrow \infty} = \infty \]

OpenStudy (dessyj1):

It is a calculus concept.

OpenStudy (dessyj1):

You would use both of them.

OpenStudy (anonymous):

okay i think i get it! thank you so much :)

OpenStudy (dessyj1):

No problem.